OpenStudy (anonymous):

Determine the number and type of solutions for 0 = x2 + 3x + 1. 1 Rational Double 2 Irrational 2 Rational 2 Complex

jimthompson5910 (jim_thompson5910):

Do you know the discriminant formula?

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

The discriminant formula is D = b^2 - 4ac

jimthompson5910 (jim_thompson5910):

In this case, a = 1, b = 3 and c = 1

jimthompson5910 (jim_thompson5910):

So D = b^2 - 4ac D = 3^2 - 4(1)(1) ... Keep going to find D